47. Which aqueous s (A) 0040 m gluo (C) 0020 m NaC 48. What is the rate law Experiment IGH 100 ml of ningen gas at the same p 44 to 46
Answer #1

42) ans = B

In given picture , He has 8 molecules.

Since nitrogen is diatomic (N2), 4 pairs of molecules must be present in the picture.

Picture I represents 4 pairs of molecules.

Therefore,

Ans = B

43)

Given that T1 = 25oC = 25 + 273 K = 298 K

V1 = 10 L

T2 = 50oC = 50 + 273 K = 323 K

V2 = ?

According to Charles law,

V1/T1 = V2/T2

10 L/298 K= V2/323K

Then, V2 = 10.8 L

Therefore,

ans = B

44) image not visible

45)

CCl4 is non-polar molecule.

The strongest intermolecular force in CCl4 is London disperson forces.

46)

  Boiling point is inversely proportional to vapor pressure.

Boiling point order is H2O > CH3OH > CH3OCH3> CH3CH3

Hence,

vapor pressure order is CH3CH3 > CH3OCH3> CH3OH> H2O

Therefore,

ans = A - I > II >IV >III